### v1

# CogSteer: Cognition-Inspired Selective Layer Intervention for Efficient Semantic Steering in Large Language Models

Abstract
Despite their impressive capabilities, large language models (LLMs) often lack interpretability and can generate toxic content. While using LLMs as foundation models and applying semantic steering methods are widely practiced, we believe that efficient methods should be based on a thorough understanding of LLM behavior. To this end, we propose using eye movement measures to interpret LLM behavior across layers. We find that LLMs exhibit patterns similar to human gaze across layers and different layers function differently. Inspired by these findings, we introduce a heuristic steering layer selection and apply it to layer intervention methods via fine-tuning and inference. Using language toxification and detoxification as test beds, we demonstrate that our proposed CogSteer methods achieve better results in terms of toxicity scores while efficiently saving 97% of the computational resources and 60% of the training time. Our model-agnostic approach can be adopted into various LLMs, contributing to their interpretability and promoting trustworthiness for safe deployment.
CogSteer: Cognition-Inspired Selective Layer Intervention for Efficient Semantic Steering in Large Language Models
[…]
1 Introduction
Large language models (LLMs) Touvron et al. (2023); Achiam et al. (2023) have demonstrated strong abilities in natural language understanding and reasoning Zhao et al. (2023). Despite the continuous increase in performance and the emergence of new capabilities from scaling LLMs Wei et al. (2022), there are significant concerns about interpretability, safety, and trustworthiness in LLMs, leading to issues such as hallucinations Huang et al. (2023) and toxicity Schick et al. (2021). As the trend of using pre-trained LLMs as foundation models, followed by fine-tuning to align with human preferences Ouyang et al. (2022); Rafailov et al. (2024), the lack of transparency in LLMs has further limited the development of efficient post-training and inference methods.
Figure 1: Framework of CogSteer. For an N-layer LLM, we first heuristically find the optimal layer M for semantic intervention. The upper block represents an adapter that is fine-tuned and inserted into the frozen layer M. The bottom block illustrates the operation of the attention module in layer M to steer the semantic direction towards safer outputs during inference.
Previous work has introduced diverse methods of interpretability, including training linear classifiers as probes on top of hidden representations Belinkov (2022), projecting representations into vocabularies Geva et al. (2022), and intervening on the computation path, such as knowledge neurons Dai et al. (2022) and circuits Conmy et al. (2023); Ghandeharioun et al. (2024). However, these methods exhibit practical shortcomings. Probings rely on limited pre-defined classes, projection methods depend on human effort to interpret limited concepts, and interventions require well-designed prompts, making them difficult to scale and generalize. Furthermore, Bills et al. (2023) uses large language models to interpret smaller language models through natural language explanations, which we find problematic, as it involves using one black box to understand another.
We believe that a more promising approach to understanding LLMs is to analyze how they operate and behave similarly to the way humans process tokens and integrate information during reading. This can be achieved by using highly interpretable indicators from eye movement studies, such as fixation and regression Rayner (1998).
In this work, we introduce a new interpretability analysis method that utilizes eye movement data Luke and Christianson (2018); Hollenstein et al. (2020); Colman et al. (2022) collected by cognition researchers to analyze human reading behaviors. We correlate the hidden states of the FFNs in each layer with distinct eye movement measures and then observe the correlation values and their changes across layers. We reveal that LLMs show patterns similar to human gaze across layers, and that different layers function differently. In particular, the middle layers, which are responsible for further token processing and information integration, are promising venues to steer semantics in LLMs.
Based on the insights gained from our interpretability analysis using eye movement measures, we propose a heuristic steering layer selection approach to identify the optimal layer for semantic intervention, which can be widely applied during fine-tuning and inference. Specifically, to assess the efficiency of our method for semantic steering, we enhance the adapter tuning method by intervening at the optimal layer instead of using the last layer or all layers by default. Additionally, we propose an implicit layer contrastive intervention method during inference, which efficiently identifies and steers semantics toward safer generation directions.
Language toxification Dementieva et al. (2024) and detoxification Leong et al. (2023) are crucial semantic steering tasks, where toxification examines the jailbreak tendencies of LLMs, and detoxification aims for safer deployment of LLMs. Experiments on these tasks demonstrate that our proposed CogSteer method, for both fine-tuning and inference, is efficient—saving 97% of computational resources and 60% of training time—and achieves better performance compared with full-layer intervention using GPT-2 (+0.9%) and LLaMa2-7B (+2.8%) models. Moreover, our proposed implicit layer contrastive intervention method not only reduces the toxicity score from 87.2% to 59.1%, but also achieves a lower score compared to the original LLaMa2-7B model at 61.0%. Lastly, these consistent results echo the findings of our interpretability analysis.
Our contributions are summarized as follows: (1) We propose a novel interpretability method that utilizes eye movement measures to perform an in-depth analysis of LLM behavior across layers. (2) Building on these insights, we introduce a heuristic steering layer selection method and apply it efficiently to semantic steering in fine-tuning and inference. (3) Through extensive experimentation and analysis, we validate the effectiveness of our proposed method on language toxification and detoxification tasks, contributing to the deployment of explainable and trustworthy foundation models.
2 Related Work
Interpretability of Large Language Models
Interpretability research is essential for uncovering the mechanisms of LLMs, ensuring their safe and trustworthy deployment as foundation models Zhao et al. (2024). Various studies have analyzed how knowledge is stored in LLMs, focusing on concepts like knowledge neurons Dai et al. (2022) and the circuits Conmy et al. (2023); Yao et al. (2024). Knowledge neurons serve as a tool for localizing specific activated neurons, while the circuits framework identifies information propagation paths Goldowsky-Dill et al. (2023); Stolfo et al. (2023). We argue that these methods are insufficient as they rely on projections into the vocabulary space Rai et al. (2024) and well-designed prompts Bills et al. (2023), using humans to interpret limited samples as concepts Geva et al. (2022). Moreover, Wang et al. (2024a) preliminarily reveals that GPT predicts tokens more similarly to humans compared with shallow language models but lacks engagement with LLMs and a broad discussion on internal mechanisms. Unlike previous works, our method explores the interpretability of LLMs through human-interpretable indicators involving eye movement theory, to precisely understand and control model behavior (see § 3.13.2).
Parameter-Efficient Semantic Steering
Parameter efficient fine-tuning (PEFT) Han et al. (2024); He et al. (2022); Wang et al. (2023) has gained popularity due to the large number of frozen parameters in LLMs for generality, while adding only a few trainable parameters per task. Adapter tuning methods Houlsby et al. (2019); Poth et al. (2023) introduce new modules between layers of LLMs, whereas prompt tuning methods Liu et al. (2022) train prompt embeddings instead. Additionally, LoRA Hu et al. (2022) improves efficiency and helps mitigate overfitting. Our method for steering semantics via fine-tuning enhances adapter tuning by incorporating a selective layer intervention strategy, saving computational resources while delivering superior performance (see § 3.3). Moreover, steering semantics during inference presents an even more efficient approach. Contrastive decoding Li et al. (2023); Sennrich et al. (2024) guides the generation direction by comparing two output distributions. Wang et al. (2024b) proposed using instruction contrast to mitigate hallucinations. In contrast, our proposed implicit layer contrastive intervention efficiently identifies and steers the semantics of LLMs toward safe directions during inference (see § 3.4).
3 Method
Figure 2: Correlation Results Under the Natural Reading Scheme. Correlation results are shown for GPT-2 Small (12 layers, top-left), GPT-2 Medium (32 layers, top-right), GPT-2 Large (36 layers, bottom-left), and LLaMa2-7B (bottom-right). Green, blue, and orange boxes indicate the premature, middle, and mature buckets, respectively.
3.1 LLMs Show Similar Pattern to Human Gaze Across Layers
Reading is a crucial ability for humans to process and integrate information, and eye movement studies enable the examination of the cognitive processes involved in reading Rayner (1998). Eye movements are monitored using various techniques, including eye-tracking systems that rely on infrared pupil monitoring. In eye-tracking studies, changes in the reader’s gaze are recorded in relation to eye movements. Similarly, tokens are processed, and information is integrated through layers in large language models Chuang et al. (2023). This similarity offers us the opportunity to leverage the extensive data previously collected and analyzed from eye movement studies, not only to infer human reading and information-processing behavior, but now to also understand the behavioral patterns of large language models.
Recent studies reveal that feed-forward networks (FFNs) function similarly to neural memory networks Geva et al. (2021), embedding syntactic features, semantic features, and even factual knowledge Chuang et al. (2023); Zhang et al. (2024). The early exit concept Elbayad et al. (2020); Schuster et al. (2022) further demonstrates that the hidden states of FFNs can be directly applied to predict words over a vocabulary. At the same time, eye movement measures provide insights into the time required for human reading to process syntax, semantics, and integrate information.
Motivated by this, we utilize eye movement measures from various datasets Luke and Christianson (2018); Hollenstein et al. (2020); Colman et al. (2022) to establish correlations with the hidden states in FFNs across layers of LLMs. Unlike traditional eye movement studies, which focus on differences in eye movement measures under experimental conditions such as natural reading or task-specific reading, we observe changes in correlation values between eye movement measures and hidden states across layers. Our approach, inspired by cognitive theory, aims to better understand the behavioral patterns of LLMs.
Formally, let $S_{j}$ denote the $j$-th sentence, which consists of $n_{j}$ words $w_{1},w_{2},\dots,w_{n_{j}}$. For each word $w_{i}$ in sentence $S_{j}$, we have five eye movement measures for each word $w_{i}$ in sentence $S_{j}$, denoted as $E_{i}^{(sfd)},E_{i}^{(ffd)},E_{i}^{(gd)},E_{i}^{(trt)},E_{i}^{(gpt)}$. Each of these represents a single scalar value per word. We also denote the hidden state at layer $l$ of the LLM for word $w_{i}$ as $\mathbf{H}_{l,i}\in\mathbb{R}^{d}$, where $d$ is the dimensionality of the hidden state vectors.
Since the eye movement measures are scalar values and each hidden state is a high-dimensional vector, we apply Principal Component Analysis (PCA) to reduce the dimensionality of the hidden states $\mathbf{H}_{l,i}$ to a scalar that can be aligned with the eye movement measure for each word $w_{i}$, yielding a one-dimensional representation: $H_{l,i}^{PCA}=PCA(\mathbf{H}_{l,i})$. We concatenate the eye movement measures and the reduced hidden states across all words in all sentences. Specifically, for the eye movement measure $k\in\{\textit{sfd},\textit{ffd},\textit{gd},\textit{trt},\textit{gpt}\}$, we concatenate the measures for all words in all sentences to form a vector $\mathbf{E}^{(k)}$:
$\mathbf{E}^{(k)}=\left[E_{1}^{(k)},E_{2}^{(k)},\dots,E_{n_{\text{total}}}^{(k)}\right]\in\mathbb{R}^{n_{\text{total}}},$ (1)
where $n_{\text{total}}=\sum_{j=1}^{m}n_{j}$ is the total number of words across all sentences. Similarly, for the PCA-reduced hidden states at layer $l$, we concatenate the reduced hidden states:
$\mathbf{H}_{l}^{\text{PCA}}=\left[H_{l,1}^{\text{PCA}},H_{l,2}^{\text{PCA}},\dots,H_{l,n_{\text{total}}}^{\text{PCA}}\right]\in\mathbb{R}^{n_{\text{total}}}.$ (2)
Finally, we compute the Pearson correlation $\rho_{l,k}$ between the hidden states at layer $l$ and the eye movement measure $k$:
$\rho_{l,k}=\frac{\sum_{i=1}^{n_{\text{total}}}\left(H_{l,i}^{\text{PCA}}-\bar{H}_{l}^{\text{PCA}}\right)\left(E_{i}^{(k)}-\bar{E}^{(k)}\right)}{\sqrt{\sum_{i=1}^{n_{\text{total}}}\left(H_{l,i}^{\text{PCA}}-\bar{H}_{l}^{\text{PCA}}\right)^{2}}\sqrt{\sum_{i=1}^{n_{\text{total}}}\left(E_{i}^{(k)}-\bar{E}^{(k)}\right)^{2}}},$ (3)
where $\bar{H}_{l}^{\text{PCA}}$ and $\bar{E}^{(k)}$ are the mean values of the reduced hidden states and the eye movement measure $k$, respectively.
We conduct a correlation analysis using the GPT-2 model at different sizes (12-layer small, 24-layer medium, 36-layer large) and the 32-layer LLaMa2-7B, following the analysis approach described earlier. To better interpret our observations, we divide the layers into three equal groups: premature, middle, and mature. Figure 2 presents the correlation results between different eye movement measures and the hidden states from the LLMs, as well as how these values change across layers.
We observe Pattern #1, where LLMs exhibit a trend similar to human gaze patterns across layers, with a peak correlation in the middle bucket. In the premature bucket, the correlation first increases, then decreases, before gradually increasing again in the later layers. In the middle bucket, the correlation reaches a peak before declining. In the mature bucket, the correlation continues to decrease, followed by a second peak. This similarity in trends is consistent across both different eye movement measures and LLMs with varying layer sizes.
Furthermore, we apply eye movement theory to interpret the phenomenon observed in Pattern #1, offering insights into the interpretability of LLMs. First fixation duration refers to the time spent on the first fixation on a word, whether it is the only fixation or the first of multiple fixations on that word. Gaze duration measures the time spent on a word during the first fixation before making a saccade to the next word. Additionally, single fixation duration refers to the time of the first and only fixation on a word.
Based on these fixation measures, the increase in correlation in the premature bucket indicates that the LLMs are beginning to process tokens by integrating syntactic and semantic features, suggesting that this bucket focuses on preliminary token processing. Geva et al. (2022) found that syntactic and semantic concepts stabilize in the first few layers using top scoring token analysis. Thus, we hypothesize that the subsequent decline in correlation may indicate that these layers are redundant but still become activated when processing more complex sentences, a point that will be explained further.
In the middle bucket, the increase in correlation signifies further processing of syntactic and semantic tokens, and the peak indicates the integration of linguistic features and information. In the mature bucket, the decline in correlation differs from that in the premature bucket. We hypothesize that the layers in the mature bucket are responsible for reasoning and factual knowledge, which will be discussed later. The second peak in the mature bucket likely represents the final integration of information for word prediction.
Figure 3: Correlation Results Comparison: Natural Reading (NR) vs. Task-Specific Reading (TSR). Correlation results are shown for GPT-2 Large NR (36 layers, top-left), GPT-2 Large TSR (36 layers, bottom-left), LLaMa2-7B NR (32 layers, top-right), and LLaMa2-7B TSR (36 layers, bottom-right). Green, blue, and orange boxes indicate the premature, middle, and mature buckets, respectively.
Moreover, total reading time represents the sum of all fixations, starting with the first fixation in a region and ending with the first forward saccade, including regressions. Go-past time refers to the total time spent on a word before moving to the right of it, including regressions, but it excludes re-fixations. Regressions are meaningful indicators, as they indicate that the reader had difficulty processing the word, failed to understand the current text, or needed to make corrections Rayner (1998).
When we zoom in on the correlation between measures like trt, gpt, and the hidden states in LLMs, the trend is consistent with the fixation-based measures. Interestingly, we observe that the correlation with trt also increases in the premature bucket. Unlike previous studies, which suggest that the first few layers of LLMs only process surface linguistic features Zhang et al. (2024), our correlation analysis shows that these layers also engage in contextual information processing.
Lastly, we compare the correlation results between GPT-2 and LLaMa2-7B to investigate the effect of modern training schemes on LLMs. The overall trend across the three buckets is similar. However, we observe three major differences. The first notable difference is that the first layer of LLaMa2-7B already shows a significant correlation with eye movement measures, likely due to the large amount of pre-training data. The second difference is the presence of a peak correlation in the middle bucket, with no evident decline afterward until reaching a second peak in the mature bucket. As discussed earlier, the premature bucket in GPT-2 is primarily responsible for token processing, while in LLaMa2-7B, it also contributes to reasoning. The last two differences will be elaborated on later. We believe that LLaMa2-7B employs modern pre-training techniques, such as in-context learning (ICL) Brown et al. (2020) and reinforcement learning with human feedback (RLHF) Ziegler et al. (2020), which distribute reasoning and factual knowledge abilities not only to the middle and mature layers but across all layers.
Empirically, LLMs are trained with the next-token prediction objective. Modern LLMs exhibit strong language understanding and reasoning abilities. This raises the question: are LLMs merely next-token predictors, or are they task reasoners? In the ZuCo 2.0 dataset Hollenstein et al. (2020), there are two different settings for eye movement studies: natural reading and task-specific reading. In the task-specific reading, participants are required to determine whether a specific relation type is present in a sentence, including categories like political affiliations, education, founder, spouse, job title, nationality, and employer. Relation detection is a high-level semantic and reasoning task that demands complex cognitive processing.
Figure 3 shows the comparison between natural reading and task-specific reading in terms of correlations between eye movement measures and hidden states in LLMs across layers. We observe Pattern #2, where LLMs function as both next-token predictors and reasoners. For the GPT-2 model, the correlation values in the premature bucket during task-specific reading are lower than those in natural reading, indicating that the layers in GPT-2’s premature bucket primarily handle token processing, not reasoning. However, in the middle bucket, the correlation peak in task-specific reading is higher than in natural reading. Furthermore, the decline in correlation across the middle and mature buckets is less pronounced, suggesting that layers in the mature bucket are particularly involved in reasoning.
For LLaMa2-7B, the correlation in the premature bucket increases in task-specific reading compared to natural reading, indicating that these layers handle both token processing and reasoning. Additionally, the correlation in both the middle and mature buckets significantly increases, suggesting that the middle bucket is responsible for token processing, information integration, and reasoning, while the mature bucket plays a key role in information integration and reasoning.
3.2 Heuristic Steering Layer Selection
A better understanding of LLM mechanisms will help in precisely and efficiently controlling their behaviors, particularly for semantic steering. We argue that the predominant parameter-efficient fine-tuning (PEFT) methods Han et al. (2024), which by default intervene in the last layer or across all layers, are not optimal [^1] Recent work Yu et al. (2023) on fine-tuning speech translation models also supports our hypothesis.. Instead, we propose an efficient heuristic steering layer selection strategy for intervention, based on our cognition-inspired interpretability analysis detailed in Section 3.1.
For semantic steering in LLMs, the layers in the middle bucket are the most suitable candidates for intervention. These layers handle further token processing, information integration, and preliminary reasoning. Additionally, the residual connections He et al. (2016) in transformer layers allow the semantic intervention to flow and evolve gradually, avoiding abrupt changes in the final prediction Chuang et al. (2023).
From a task-oriented perspective, we apply PEFT methods, such as Adapters Poth et al. (2023), or inference-only methods to the candidate layers in the middle bucket, using a small portion of data, like the validation set, to search for and select the best-performing layer that suits the task scenario. Formally, given $M$ as the best layer for intervention, $J$ represents a set of candidate layers in the middle bucket $\left(\frac{N}{3}\leq M\leq\frac{2N}{3}\right)$. $D$ denotes the validation set. We search for the layer $M$ that yields the best task score or loss performance, as follows:
$M=\arg\max_{j\in J}S\left(D;P\left(\cdot\mid x_{t}\right)\right).$ (4)
Later, we will demonstrate the use of the heuristic steering layer selection approach in the context of language toxification and detoxification tasks, as elaborated in Section 4.
3.3 Layer Intervention via Fine-tuning
Adapter methods are among the most popular and widely used PEFT approaches due to their efficiency, requiring only a small set of new parameters for task-specific fine-tuning. Let the parameters of an LLM consist of a set of pre-trained, frozen parameters $\phi(\cdot)$ and a set of newly introduced parameters in the adapter block $\psi(\cdot)$. Our layer intervention via fine-tuning to steer semantics in LLM and predicts the next token as follows:
$y\left(x_{t}\right)=\operatorname{softmax}\left(\operatorname{logit}_{\phi,\psi}\left(FFN_{\phi}^{N}\left(x_{t}\right)\mid F_{\phi,\psi}^{M}\left(x_{t}\right),y_{<t}\right)\right).$ (5)
Here, $FFN_{\phi}^{N}(x_{t})$ represents the hidden state of the FFN in the final layer with the frozen parameters $\phi(\cdot)$, used for token prediction over a vocabulary. $M$ denotes the best layer for semantic intervention, determined by Equation 4. $F_{\phi,\psi}^{M}(x_{t})$ indicates that the fusion layer in the LLM incorporates new parameters $\psi(\cdot)$ from the adapter block, aligned with the frozen parameters $\phi(\cdot)$.
Our cognitive-inspired selective layer intervention method is an adaptive fine-tuning strategy that identifies the best layer for both effective semantic steering and task performance. Moreover, as our method only operates on a single layer rather than all layers, it significantly reduces computational resources and time, while also avoiding catastrophic forgetting Luo et al. (2023); Li et al. (2024).
3.4 Layer Intervention during Inference
Efficient semantic steering can be achieved via fine-tuning. However, an even more efficient approach is to steer the semantics of LLMs during inference without introducing additional parameters. Motivated by Li et al. (2023); Leong et al. (2023), which contrast outputs from either a less capable model or outputs induced from a negative prompt, we propose an implicit layer contrastive intervention method during inference. First, we fine-tune a contrast model that generates either the desired output or the output we aim to mitigate. In our case, to mitigate toxic token generation, we fine-tune a toxic LLM as the contrast model. Unlike Li et al. (2023); Leong et al. (2023), which contrast the outputs explicitly in the last layer, our method operates on the contextualized value vectors derived from the weight matrices $K$, $Q$, $V$ of the attention modules within LLMs (see Elhage et al. (2021) for mathematical derivation). We perform this operation on the best layer for intervention as described in Equation 4. Formally, our layer intervention during inference finds the semantic steering direction by contrasting the value vectors as follows:
$\Delta v^{M}=v_{c}^{M}-v_{o}^{M},$ (6)
where $v_{c}^{M}$ and $v_{o}^{M}$ are the contextualized value vectors of the contrast LLM and the original LLM at the best layer $M$ for semantic intervention. We then update the value vector in the layer $M$ of the original LLM:
$v^{\prime M}=v_{o}^{M}-\lambda_{\text{norm}}^{\alpha}\cdot\Delta v^{M}.$ (7)
Here, $\lambda_{\text{norm }}=1+\left\|\Delta v^{M}\right\|_{2}$ is a normalization term that adaptively affects the steering effect, and $\alpha$ is a hyperparameter that further controls the steering strength. Finally, we keep the updated steering direction and renormalize the adapted value vector to ensure its representation is close to the original vector:
$v^{\prime M}=v^{\prime M}\cdot\frac{\left\|v_{c}^{M}\right\|_{2}}{\left\|v^{\prime M}\right\|_{2}}.$ (8)
Figure 4: Toxification/Detoxification Evaluation. The bottom chart shows toxicity scores for interventions at each layer: green bars represent selective intervention via fine-tuning, and purple bars represent selective intervention during inference. The top chart displays detoxify margin scores per layer, comparing toxicity scores in toxification and detoxification processes. Purple, blue, and orange boxes indicate the premature, middle, and mature buckets.
4 Experiments
4.1 Datasets and Evaluation
Datasets We consider natural language toxification and detoxification tasks as dual tasks to examine our proposed efficient semantic steering methods through selective layer intervention. To train toxic adapters and contrast models, as described in Sections 3.3 and 3.4, we use the Toxic Comment Classification Challenge Dataset Jigsaw (2018), which contains 15,294 annotated toxic comments. We randomly split the dataset into 13,764 comments for fine-tuning and training the contrast models, and 1,530 comments for the validation set, aimed at identifying the best layer for intervention as described in Section 3.2. For evaluation, we use the RealToxicityPrompts (RTP) dataset Gehman et al. (2020). Following the setup of prior work Leong et al. (2023), we sample 2,122 toxic prompts to assess the performance of our proposed CogSteer method during fine-tuning and inference.
Evaluation For each toxic prompt in the RTP dataset, we generate 25 continuations, following the setup in Li et al. (2023); Leong et al. (2023). We then evaluate the toxicity of these continuations using the Perspective API  [^2] https://perspectiveapi.com/, which assigns a toxicity score to each one. This score represents the probability that a continuation is perceived as toxic, with higher scores indicating a greater likelihood of toxicity. Finally, we use the average maximum toxicity as our evaluation metric. This metric involves identifying the highest toxicity score from the 25 continuations for each prompt, and then calculating the average of these maximum scores across all prompts.
4.2 Models and Baselines
We evaluate our efficient semantic steering methods using three sizes of GPT-2 models (12-layer small, 24-layer medium, and 36-layer large) and LLaMa2-7B. We selected the GPT-2 model and LLaMa2-7B as the former represents earlier classical LLMs, while the latter represents modern LLMs with advanced techniques such as ICL and RHLF. For generalizability, we believe our proposed method is LLM-agnostic and can be applied to other LLMs. For the Adapter block in GPT-2 models, we used the adapter implementation Poth et al. (2023) from HuggingFace, while for LLaMa2-7B we used the LLaMa-Adapter implementation Zhang et al. (2023). Our experiments focus on the performance of selective layer intervention. Specifically, we compare the toxicity scores and detoxify margins when applying our methods to different layers of the models, with the best intervention layer selected under the same task. Since applying semantic intervention to the last layer and all layers of LLMs is a conventional approach, we use these two methods as baselines for comparison.
4.3 Analysis on Language Toxification
We evaluate our selective layer intervention via fine-tuning using the language toxification task. We generate continuations for each toxic prompt in the RTP dataset and compute and compare the average maximum toxicity. Figure 4 (green bars) shows the toxicity score for inserting a toxic adapter into each layer of the LLMs, with the last layer and full layers (red line) used as baselines. The best layer $M$ in the middle bucket for semantic intervention is $L23$ for GPT-2 and $L13$ for LLaMa2-7B, competing with the results from both the last layer (+4.5% for GPT & +23% for LLaMa) and full layers (+0.9% for GPT & +2.8% for LLaMa). This indicates that the conventional intervention in the last layer or full layers is not optimal, demonstrating the effectiveness of our proposed heuristic steering layer selection approach, which saves considerable computational resources and time for fine-tuning.
Moreover, for LLaMa2-7B, the toxicity scores for layer intervention in the middle bucket $L11$-$L17$ are significant compared to both the last layer and full-layer interventions. In contrast, for GPT-2, the toxicity scores for layer intervention in $L21$-$L24$ in the middle bucket are competitive with last layer and full-layer interventions, showing similar performance. These strong performances for layer-wise semantic intervention align with our analysis based on eye movement measures, which shows that the middle bucket layers are involved in further token processing and information integration.
Furthermore, layer intervention in the mature bucket, including the last layer (used as a baseline), is not remarkable when compared to full-layer intervention for both GPT-2 and LLaMa2-7B models. According to our behavioral analysis in Section 3.1, the layers in the mature bucket focus more on reasoning and factual knowledge, involving less token processing, which makes them less suitable for semantic steering. Interestingly, for the GPT-2 model, we observe that layers $L1$-$L4$ and $L6$-$L8$ in the premature bucket also outperform the full-layer intervention baseline, whereas this phenomenon is not observed in LLaMa2-7B. We hypothesize that this difference may be due to the distinct adapter training schemes. For GPT-2, we use the vanilla Adapter as a plug-in, which directly affects token distributions and allows successive minor changes as tokens pass through layers, following the early exit theory Elbayad et al. (2020); Schuster et al. (2022). However, for LLaMa2, we use the LLaMa-Adapter, which directly modifies attention. Edits to the attention modules gradually affect token distributions Geva et al. (2021); Elhage et al. (2021).
4.4 Analysis on Language Detoxification
As a dual task, we evaluate our selective layer intervention during inference using the language detoxification task. This task serves as an adversarial task to mitigate the toxic tokens that are amplified by layer intervention methods during fine-tuning. Thus, we observe the detoxify margin in each intervention layer and compare the performance with intervention in the last layer. Figure 4 (blue bars in the upper chart) shows the results of the detoxify margin scores. The best layer $M$ for semantic intervention in the middle bucket is $L23$ for GPT-2 and $L16$ for LLaMa2-7B, both outperforming the last layer results by a large margin (+2.9% for GPT and +24% for LLaMa). Notably, the best layer $M$ for LLaMa2-7B in the toxification task is $L13$, whereas the best layer for detoxification is $L16$, demonstrating that our heuristic layer selection method is adaptive and suitable for tasks.
Moreover, the detoxify margin scores for layer intervention in the middle bucket for both GPT-2 and LLaMa2 models are significantly better than those for intervention in the mature bucket. This aligns with our findings that the middle bucket layers are involved in further token processing and information integration, whereas the mature bucket layers focus on reasoning, making them less suitable for semantic steering. Finally, for the GPT-2 model, the detoxify margin scores are also significant in the premature bucket layers. This could be attributed to the distinct adapter training scheme, as the earlier layers are more deeply affected by the toxification, which broadens the scope of detoxification. The remaining results for GPT-2 small and medium models are listed in Appendix B.
4.5 Efficiency Analysis
The efficiency analysis of our proposed cognition-inspired selective layer intervention is presented in Table 1. By selecting the best layer for semantic steering in LLMs, the training time and number of parameters involved in fine-tuning are significantly reduced compared to full-layer intervention. Specifically, our method requires only 40% of the time and 3% of the parameters needed for full-layer settings. Importantly, this reduction in computational cost does not compromise performance; in fact, layer intervention is comparable to, and in some cases even better than, full-layer intervention across all models.
[…]
GPT2-S → L7 5 min 0.42 0.15 M $0.88_{-0.02}$ $0.60_{-0.01}$
GPT2-S → Full 12 min 1.7 M 0.90 0.59
GPT2-M → L14 9 min 0.38 0.26 M $0.85_{+0.02}$ $0.63_{+0.01}$
GPT2-M → Full 24 min 6.3 M 0.87 0.64
GPT2-L → L22 13 min 0.39 0.41 M $0.87_{+0.01}$ $0.63_{-0.03}$
GPT2-L → Full 33 min 14 M 0.86 0.60
Llama2-7B → L15 85 min 0.41 0.04 M $0.87_{+0.01}$ $0.59_{+0.03}$
Llama2-7B → Full 205 min 1.3 M 0.86 0.62
Table 1: Efficiency Comparison between CogSteer and Full-Layer Intervention
5 Conclusion
In this paper, we introduce an efficient semantic steering method for LLMs using cognition-inspired selective layer intervention. Our approach is inspired by correlation analysis with eye movement measures, making our method understandable and highly interpretable. Specifically, we select the optimal layer for intervention based on heuristic steering layer selection and then apply it to efficient steering methods during fine-tuning and inference. Experimental results show that CogSteer significantly mitigates the toxicity of LLMs while being efficient in terms of training time and parameters involved. Overall, our proposed method is an important step toward making LLMs more interpretable and contributes to the safe deployment of LLMs.
Limitations
In this paper, we conduct a thorough analysis of the correlation between the hidden states in the feed-forward network and eye movement measures. However, there is another key component in Transformers: the attention block. We limit our discussion to the FFN blocks because the hidden states in FFNs are more responsible for token prediction. For future work, we will explore the mechanisms of the attention block in greater depth. Additionally, we conclude that the layers in the mature buckets are involved in reasoning based on the task-specific reading dataset. However, factual knowledge also seems to be integrated in this bucket. We will seek further eye movement studies that involve factual information processing to refine our analysis.
Ethics Statement
We propose an efficient semantic steering method in LLMs through cognition-inspired selective layer intervention to better understand the behavior of LLMs and address safety concerns, thereby enhancing their safety and reliability within the community. Additionally, the eye-tracking data used in this study, derived from the ZuCo 2.0, GeCo, and Provo datasets, are publicly available and adhere to established ethical protocols, promoting transparency and reproducibility in our research. Furthermore, we have made our code publicly accessible, ensuring that researchers and practitioners can easily access and implement our methods.
Acknowledgments
This research was funded by the German Research Foundation DFG Transregio SFB 169: Crossmodal Learning: Adaptivity, Prediction, and Interaction.
Appendix A Implementation Details
For training the GPT-2 adapters, we set the learning rate to $5\times 10^{-4}$ and train for 5 epochs. For the LLaMA2-7B adapters, we use the default settings provided by LLaMA-Adapter Zhang et al. (2023). In all models, we apply the implicit layer contrastive intervention approach with $\alpha=0.4$. Following previous works Liu et al. (2021); Leong et al. (2023), the model generates 25 continuations per prompt using nucleus sampling with $p=0.9$, with each continuation limited to a maximum of 20 tokens.
Appendix B Further Analysis on Language Toxification and Detoxification
Figure 5: Toxification/Detoxification Evaluation on GPT-2 Small (12 Layers). The bottom chart shows toxicity scores for interventions at each layer: green bars represent selective intervention via fine-tuning, and purple bars represent selective intervention during inference. The top chart displays detoxify margin scores per layer, comparing toxicity scores in toxification and detoxification processes. Purple, blue, and orange boxes indicate the premature, middle, and mature buckets.
Figure 6: Toxification/Detoxification Evaluation on GPT-2 Medium (24 Layers). The bottom chart shows toxicity scores for interventions at each layer: green bars represent selective intervention via fine-tuning, and purple bars represent selective intervention during inference. The top chart displays detoxify margin scores per layer, comparing toxicity scores in toxification and detoxification processes. Purple, blue, and orange boxes indicate the premature, middle, and mature buckets.
Appendix C Qualitative Analysis
Warning: Some examples have harmful or offensive language.
Figure 7: Case Study